This is a differential equations course, please answer all the questions.

Find the general solution.

12. z(t) -1)=(0) 13. z(t) = -3)=(0) 14. zº(t) = (+5 -1)=(e) 5 10 0 15. x(t) = 21 -2 r(t) 3 2 -1) 16. z(t) = -30 2 1 -1 0
Answer #1

SOLUTION 72) (t) = 2. I Mih nino X(t) r(t) = AX(+) Eigen value equation for Coefficient matrix /A-AT/=0 Q- --- ule -/- (2-xR2 R2 - RI 5(3-5) 3-85 2 - 2 7[*]+[:] Here 22 is free variable Choose X2 = 1 3-5 2 x 24 Ş X2 = 0 2 / 2 3-85 5 ( 3+15) (3-15)3+ √5 - 24 [20 -3+VE 2 R2 R2- 5 (3+15) STAR so :] [2][ ~ 3+v5 2 Here X is free variable choose X2=1 3+05 2 21 E x2 = 0 Z 5 3+Hence complele Solution: mit + cx xr ett X (t) = ex, e za se (1+5)* X (t) { (3+V5) 7 e (-V5 ) + **(****) 2170

Solved only 12)